Kate is 5 years younger than Susan, and Melinda is 2 years older than Susan. If the sum of all three of their ages is 36, how old is Melinda?
8

Answer by math_helper(2461)

K = S - 5
S = S
M = S + 2
K+S+M = (S-5)+S+(S+2) = 36
3S = 39
S = 13, K = 8, and M = 15